OpenStudy (anonymous):

help @alexandervonhumboldt2

OpenStudy (anonymous):

The table shows data for a class's mid-term and final exams: Mid-Term Final 100 98 100 95 100 93 95 91 95 88 92 82 92 78 88 78 85 65 75 60 Which data set has the largest IQR?

OpenStudy (anonymous):

@texaschic101 @tkhunny @timo86m @TheRaggedyDoctor @unknown101 @amistre64

OpenStudy (tkhunny):

You should calculate it. What are the medians of the two data sets?

OpenStudy (anonymous):

I really forgot

OpenStudy (tkhunny):

It's time to remember. You need a value with equal numbers of data on each side. Go!

OpenStudy (anonymous):

set 3???

OpenStudy (tkhunny):

What is the median of this set of data: 1 3 7

OpenStudy (anonymous):

um...7

OpenStudy (tkhunny):

Does 7 have an equal number of data items on each side?

OpenStudy (anonymous):

no.

OpenStudy (tkhunny):

Then it is not 7. What IS the median?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

median is middle right??

OpenStudy (tkhunny):

That is right. It has one data item on each side. Tougher question. What is the median of this data set: 1 3 5 7

OpenStudy (anonymous):

subtract 5 and 3???

OpenStudy (tkhunny):

Not bad. The only way to get and equal number on each side is to pick something between 3 and 5. Usually, that's (3+5)/2 = 8/2 = 4 Now to your data... What is the median of this set? 100 100 100 95 95 92 92 88 85 75

OpenStudy (anonymous):

93

OpenStudy (anonymous):

so it is set 3

OpenStudy (tkhunny):

That makes no sense. There are only two sets. 93 is a good median for the mid term. What is a good median for the final?

OpenStudy (tkhunny):

gtg Once you find that median, you can use those two medians to think of both sets being divided into two subsets each. Find the median of all four subsets and you will be ready to calculate the IQR.